+ 0/1.2 points | Previous Answers Tipler6 24.P.045 My Notes Three concentric, thin, long conducting cylindrical shells have radii of 2.00 mm, 6.00 mm, and 7.00 mm. The space between the shells is filled with air. The innermost and outermost cylinders are connected at one end by a conducting wire. Find the capacitance per unit length of this configuration. 17.9 pF/m eBook Submit Answer Save Progress
Answer #1

here it behaves like two capacitors connected in parallel.


let a = 2 mm, b = 6 mm, c = 7 mm

C1/L = 2*pi*epsilon/ln(b/a)

= 2*pi*8.854*10^-12/ln(6/2)

= 5.06*10^-11 F/m

C2/L = 2*pi*epsilon/ln(c/b)

= 2*pi*8.854*10^-12/ln(7/6)

= 3.61*10^-10 F/m


Ceq/L = C1/L + C2/L

= 5.06*10^-11 + 3.61*10^-10

= 4.12*10^-10 F/m

= 412*10^-12 F/m

= 412 pF/m <<<<<<<--------------Answer